# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ 3.2 Metric properties of TpOT
It is a classic result that the coupling space $\Pi(\mu,\mu^{\prime})$ is (sequentially) compact in $P(X\times X^{\prime})$, where $P(X\times X^{\prime})$ is the set of Borel probability measures on $X\times X^{\prime}$, see, e.g.[50, Lemma 4.4] and [11, proof of Lemma 24], where we topologise spaces of measures with the weak topology. Similarly, $\Pi_{\rm adm}(\nu,\nu^{\prime})$ is sequentially compact in the space $M(\overline{Y}\times\overline{Y^{\prime}})$ of Radon measures on $\overline{Y}\times\overline{Y^{\prime}}$ (see [17, proof of Prop 3.1], which extends to Polish spaces as in Definition 1), with the vague topology. These two together imply that $\Pi(\mu,\mu^{\prime})\times\Pi_{\rm adm}(\nu,\nu^{\prime})$ is sequentially compact in $P(X\times X^{\prime})\times M(\overline{Y}\times\overline{Y^{\prime}})$ with the product topology.
@@
$C_{p}(\pi,\xi)=\ \int_{X\times X^{\prime}\times\overline{Y}\times\overline{Y^{\prime}}}|\omega-\omega^{\prime}|^{p}\>\mathrm{d}(\pi\otimes\xi)+\int_{(X\times X^{\prime})^{2}}|k-k^{\prime}|^{p}\>\mathrm{d}(\pi\otimes\pi)+\int_{\overline{Y}\times\overline{Y^{\prime}}}|d_{p}\circ(\iota,\iota^{\prime})|^{p}\>\mathrm{d}\xi.$ (10)
@@
To prove Prop. 2 we only need to show that the triangle inequality holds, as symmetry is obvious from the definition of $d_{\mathrm{TpOT},p}$. We follow the usual approach, which relies on the following standard result (see e.g. [43, Lemma 1.4]).
@@
Let $(X_{i},\mu_{i}),i=1,\cdots,n$, be Polish probability spaces. Consider couplings $\xi_{i}\in\Pi(\mu_{i-1}\mu_{i})$, for $i=2,\cdots,n$. There exists a unique probability measure $\xi$ on $X_{1}\times\cdots\times X_{n}$ such that $(p_{i-1}\times p_{i})_{\#}\xi=\xi_{i}$ for every $i=2,\cdots,n$, where $p_{i}:X_{1}\times\cdots\times X_{n}\longrightarrow X_{i}$ is the projection on the $i$th factor.
The unique measure guaranteed by the lemma will be referred to as the gluing of $\xi_{2},\ldots,\xi_{n}$ and will sometimes be denoted as $\xi_{2}\boxtimes\xi_{3}\boxtimes\cdots\boxtimes\xi_{n}$. Lemma 1 can be adapted to the case of Radon measures and admissible couplings [17, Prop. 3.2].
@@
The proof follows adapts techniques used in the proof of [17, Prop. 3.2] and [11, Thm. 1]. To prove that the triangle inequality holds for $d_{\mathrm{TpOT},p}$, consider three topological networks $P_{1}$, $P_{2}$ and $P_{3}$, and let $(\pi_{12}^{v},\pi^{e}_{12})$ and $(\pi_{23}^{v},\pi^{e}_{23})$ be the couplings realising $d_{\mathrm{TpOT},p}(P_{1},P_{2})$ and $d_{\mathrm{TpOT},p}(P_{2},P_{3})$, which exist by Prop. 1. Then, using the gluing lemma, we construct a probability measure $\pi^{v}$ on $X_{1}\times X_{2}\times X_{3}$ with marginals $\pi^{v}_{12}$ and $\pi^{v}_{23}$ on $X_{1}\times X_{2}$ and $X_{2}\times X_{3}$, respectively. We denote by $\pi^{v}_{13}$ the marginal on $X_{1}\times X_{3}$. Similarly, the gluing lemma adapted to Radon measures and admissible couplings yields a measure $\pi^{e}$ on $\overline{Y_{1}}\times\overline{Y_{2}}\times\overline{Y_{3}}$ with marginals that agree with $\pi^{e}_{ij}$ when restricted to $\overline{Y_{i}}\times\overline{Y_{j}}\setminus\{(\partial_{Y_{i}},\partial_{Y_{j}})\}$ and that induces a zero cost on $(\partial_{Y_{i}},\partial_{Y_{j}})$. Then, we have
$d_{\mathrm{TpOT},p}(P_{1},P_{3})$ $\leq\left(\|d_{p}\circ(\iota_{1},\iota_{3})\|^{p}_{L^{p}(\pi^{e}_{13})}+\|k_{1}-k_{3}\|^{p}_{L^{p}(\pi^{v}_{13}\otimes\pi^{v}_{13})}+\|\omega_{1}-\omega_{3}\|^{p}_{L^{p}(\pi^{v}_{13}\otimes\pi^{e}_{13})}\right)^{1/p}$ (11)
@@
Here, we use $\|\cdot\|_{p}$ to indicate the standard $\ell^{p}$ norm on $\mathbb{R}^{3}$. The inequality (11) follows by suboptimality, while (12) and (15) by marginal conditions. The triangle inequalities in $L^{p}$ and $d_{p}$ give (13), and the triangle inequality in $\ell^{p}$ gives (14). ∎
@@
Two topological measure networks $P=\left((X,k,\mu),(Y,\iota,\nu),\omega\right)$ and $P^{\prime}=\left((X^{\prime},k^{\prime},\mu^{\prime}),(Y^{\prime},\iota^{\prime},\nu^{\prime}),\omega^{\prime}\right)$ have distance $d_{\mathrm{TpOT},p}(P,P^{\prime})=0$ if and only if $P\sim_{w}P^{\prime}$.
@@
The “if” direction is clear, by the triangle inequality. Conversely, assume $d_{\mathrm{TpOT},p}(P,P^{\prime})=0$ and let $\pi^{v}$ and $\pi^{e}$ be couplings realising this distance (Prop. 1).
@@ 3.3 A topology-driven metric on point clouds
We now summarise the metric described above in the discrete setting of point clouds with finite number of points. Consider a point cloud $X=\{x_{1},\cdots,x_{N}\}$ in $\mathbb{R}^{n}$. Any choice of filtration $\mathcal{K}$ over $X$ yields a corresponding persistence diagram $D=D(X,\mathcal{K})$. Choosing a representative cycle for each homology class in $D$ yields the set of generators $g$, and the corresponding PH-hypergraph $H=H(X,\mathcal{K},g)$. We can now endow $X$ with a topological measure hypergraph structure by considering
@@
Given two point clouds $X=\{x_{1},\cdots,x_{N}\}$ and $X^{\prime}=\{x^{\prime}_{1},\cdots,x^{\prime}_{N^{\prime}}\}$ in $\mathbb{R}^{n}$ endowed with filtrations, let $P_{X}$ and $P_{X^{\prime}}$ be their associated topological networks. In this concrete setting, we focus on the $p=2$ version of the metric. As was described in Remark 2, it is useful in applications to include tunable weights on the terms of TpOT. With this notation, the distance $d_{\mathrm{TpOT},2}=d_{\mathrm{TpOT},\alpha,\beta}$ can be reformulated, using notation similar to that of [10], as
$d_{\mathrm{TpOT},\alpha,\beta}(P_{X},P_{X^{\prime}})^{2}=\min_{\pi^{v}\in\Pi(\mu,\mu^{\prime}),\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\alpha\left\langle L(C,C^{\prime}),\pi^{v}\otimes\pi^{v}\right\rangle+(1-\alpha)\left\langle\widetilde{C}(D,D^{\prime}),\pi^{e}\right\rangle+\beta\left\langle L(\omega_{H},\omega_{H^{\prime}}),\pi^{v}\otimes\pi^{e}\right\rangle.$ (16)
@@
  • All inner products in this expression are Frobenius inner products of matrices.
  • The terms $C,C^{\prime}$ are the pairwise (Euclidean) distance matrices, and $D,D^{\prime}$ are the persistence diagrams associated to the given filtrations.
  • The term $\widetilde{C}(D,D^{\prime})$ denotes the augmented $L^{2}$ cost matrix of dimension $(|D|+1)\times(|D^{\prime}|+1)$ for persistence diagrams, whose last column and row correspond to transport to and from $\partial_{\Lambda}$ respectively, following the diagonal projection $\pi_{\partial_{\Lambda}}$ (see [28, Equation 8]).
  • The expression $L(C,C^{\prime})$ denotes the fourth-order distortion tensor, whose term with index $(i,j,k,\ell)$ is given by $L(C,C^{\prime})_{ijk\ell}=\frac{1}{2}|C_{ik}-C^{\prime}_{j\ell}|^{2}$. The term $\langle L(C,C^{\prime}),\pi^{v}\otimes\pi^{v}\rangle$ can be understood by considering the measure $\pi^{v}$ as a matrix, so that the product $\pi^{v}\otimes\pi^{v}$ is also naturally identified with a fourth order tensor (with the same dimensions as $L(C,C^{\prime})$). Then the inner product is computed by isomorphically identifying these fourth order tensors with matrices.
  • We denote by $\omega_{H}$ and $\omega_{H^{\prime}}$ the hypernetwork functions of the PH-hypergraphs $H$ and $H^{\prime}$, respectively. Specifically, $\omega_{H}$ and $\omega_{H^{\prime}}$ are represented as matrices of size $N\times|D|+1$ , $N^{\prime}\times|D^{\prime}|+1$ respectively, that are obtained as the hypergraph incidence matrices with an additional 00-row each representing membership for the diagonal edge.
  • The expression $L(\omega_{H},\omega_{H^{\prime}})$ is the fourth-order given by:
[…]
    $\big{(}L(\omega_{H},\omega_{H^{\prime}})\big{)}_{|D|+1,k,|D^{\prime}|+1,l}$ $=0$ (19)
    In other words, the cost of matching any real edge to the diagonal is given by the $L^{2}$ norm of the edge (17 and 17), while matching the diagonal with the diagonal has zero cost (19).
@@
The solution of the TpOT problem in 16 involves determining an optimal pair of couplings $\pi^{v},\pi^{e}$. The coupling $\pi^{v}$ induces a transport plan between points in the point clouds $X,X^{\prime}$. When $\beta>0,\alpha\in[0,1]$, this transport balances between preserving topological features and pairwise distances. The coupling $\pi^{e}$ induces a matching between homology classes in the persistence diagrams $D,D^{\prime}$. For $10<\alpha<$ 1, this matching is driven by the Wasserstein matching, and it is informed by pairwise Euclidean proximity of points forming representative cycles of classes in $D,D^{\prime}$. Implementation details are described in Section 5, where we also provide numerical examples.
@@ 3.4 Numerical algorithms
We now aim at numerical algorithms for approximating the solution to the TpOT problem in practice. Starting from (16), we consider a regularised variant of the TpOT problem by adding an entropic regularisation to the transport plans $(\pi^{e},\pi^{v})$. Writing
$\mathsf{L}(\pi^{v},\pi^{e})=\alpha\langle L(C,C^{\prime}),\pi^{v}\otimes\pi^{v}\rangle+(1-\alpha)\langle\widetilde{C}(D,D^{\prime}),\pi^{e}\rangle+\beta\langle L(\omega_{H},\omega_{H^{\prime}}),\pi^{v}\otimes\pi^{e}\rangle,$
@@
$\min_{\pi^{v}\in\Pi(\mu,\mu^{\prime}),\>\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}$ $\mathsf{L}(\pi^{v},\pi^{e})+\varepsilon_{v}\operatorname{KL}(\pi^{v}|\mu\otimes\mu^{\prime})+\varepsilon_{e}\operatorname{KL}(\pi^{e}|\widetilde{\nu}\otimes\widetilde{\nu}^{\prime}).$
For $\varepsilon_{v},\varepsilon_{e}>0$, we can take advantage of the entropic regularisation to utilise fast, smooth optimisation techniques. In particular, we can find a local minimum by projected gradient descent [37, Section 2.3], where both the gradient and the projection are calculated with respect to the KL-divergence. This leads to the following iterative scheme:
$\begin{split}\pi^{v}_{t+1}&\leftarrow\operatorname{Proj}_{\Pi(\mu,\mu^{\prime})}^{\operatorname{KL}}\left[-\varepsilon_{v}^{-1}\nabla_{v}\mathsf{L}(\pi^{v}_{t},\pi^{e}_{t})(\mu\otimes\mu^{\prime})\right]\\ \pi^{e}_{t+1}&\leftarrow\operatorname{Proj}_{\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}^{\operatorname{KL}}\left[-\varepsilon_{e}^{-1}\nabla_{e}\mathsf{L}(\pi^{v}_{t},\pi^{e}_{t})(\widetilde{\nu}\otimes\widetilde{\nu}^{\prime})\right].\end{split}$ (20)
Each of these projections can be calculated by matrix scaling using the Sinkhorn algorithm [36, Chapter 4], and the gradients of $\mathsf{L}$ have closed-form expressions:
$\begin{split}\nabla_{v}\mathsf{L}(\pi^{v},\pi^{e})&=2\alpha L(C,C^{\prime})\otimes\pi^{v}+\beta L(\omega_{H},\omega_{H}^{\prime})\otimes\pi^{e}\\ \nabla_{e}\mathsf{L}(\pi^{v},\pi^{e})&=(1-\alpha)\widetilde{C}[D,D^{\prime}]+\beta L(\omega_{H}^{\top},{\omega_{H}^{\prime}}^{\top})\otimes\pi^{v}.\end{split}$ (21)
In the unregularised case when $\varepsilon_{v}=\varepsilon_{e}=0$, an alternating minimisation in $(\pi^{v},\pi^{e})$ can be used to find a local minimum. Fixing $\pi^{e}$, the minimisation problem in $\pi^{v}\in\Pi(\mu,\mu^{\prime})$ is
@@
This falls within the fused Gromov-Wasserstein framework which was introduced and studied in detail by [45, 48]. A local minimum can be found using a conditional gradient method [45, Algorithm 1].
Fixing $\pi^{v}$ and minimising in $\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})$, we have
$\min_{\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\>\langle M,\pi^{e}\rangle,\quad M=\pi^{v}\otimes L(\omega_{H},\omega_{H^{\prime}})+(1-\alpha)\widetilde{C}(D,D^{\prime})$ (23)
and we have the identity
@@
This amounts to a standard optimal transport problem.

### v2

@@ 3.2 Metric properties of TpOT
It is a classic result that the coupling space $\Pi(\mu,\mu^{\prime})$ is (sequentially) compact in $P(X\times X^{\prime})$, where $P(X\times X^{\prime})$ is the set of Borel probability measures on $X\times X^{\prime}$, see, e.g.[61, Lemma 4.4] and [13, proof of Lemma 24], where we topologise spaces of measures with the weak topology. Similarly, $\Pi_{\rm adm}(\nu,\nu^{\prime})$ is sequentially compact in the space $M(\overline{Y}\times\overline{Y^{\prime}})$ of Radon measures on $\overline{Y}\times\overline{Y^{\prime}}$ (see [19, proof of Prop 3.1], which extends to Polish spaces as in Definition 1) with the vague topology, i.e. given by duality between measures and continuous functions with compact support. Together, these imply that $\Pi(\mu,\mu^{\prime})\times\Pi_{\rm adm}(\nu,\nu^{\prime})$ is sequentially compact in $P(X\times X^{\prime})\times M(\overline{Y}\times\overline{Y^{\prime}})$ with the product topology.
@@
$\begin{split}C_{p}(\pi,\xi)&=\ \int_{X\times X^{\prime}\times\overline{Y}\times\overline{Y^{\prime}}}|\omega-\omega^{\prime}|^{p}\>\mathrm{d}(\pi\otimes\xi)\\ &\qquad\qquad+\int_{(X\times X^{\prime})^{2}}|k-k^{\prime}|^{p}\>\mathrm{d}(\pi\otimes\pi)+\int_{\overline{Y}\times\overline{Y^{\prime}}}|d_{p}\circ(\iota,\iota^{\prime})|^{p}\>\mathrm{d}\xi.\end{split}$ (11)
@@
To prove Proposition 2 we only need to show that the triangle inequality holds, as symmetry is obvious from the definition of $d_{\mathrm{TpOT},p}$. We follow the usual approach, which relies on the following standard result (see e.g. [54, Lemma 1.4]).
@@
Let $(X_{i},\mu_{i}),i=1,\cdots,n$, be Polish probability spaces. Consider couplings $\xi_{i}\in\Pi(\mu_{i-1},\mu_{i})$, for $i=2,\cdots,n$. There exists a unique probability measure $\xi$ on $X_{1}\times\cdots\times X_{n}$ such that $(p_{i-1}\times p_{i})_{\#}\xi=\xi_{i}$ for every $i=2,\cdots,n$, where $p_{i}:X_{1}\times\cdots\times X_{n}\longrightarrow X_{i}$ is the projection on the $i$th factor.
The unique measure guaranteed by the lemma will be referred to as the gluing of $\xi_{2},\ldots,\xi_{n}$ and will sometimes be denoted as $\xi_{2}\boxtimes\xi_{3}\boxtimes\cdots\boxtimes\xi_{n}$. Lemma 1 can be adapted to the case of Radon measures and admissible couplings [19, Proposition 3.2].
@@
The proof follows adapts techniques used in the proof of [19, Proposition 3.2] and [13, Theorem 1]. To prove that the triangle inequality holds for $d_{\mathrm{TpOT},p}$, consider three topological networks $P_{1}$, $P_{2}$ and $P_{3}$, and let $(\pi_{12}^{v},\pi^{e}_{12})$ and $(\pi_{23}^{v},\pi^{e}_{23})$ be the couplings realising $d_{\mathrm{TpOT},p}(P_{1},P_{2})$ and $d_{\mathrm{TpOT},p}(P_{2},P_{3})$, which exist by Prop. 1. Then, using the gluing lemma, we construct a probability measure $\pi^{v}$ on $X_{1}\times X_{2}\times X_{3}$ with marginals $\pi^{v}_{12}$ and $\pi^{v}_{23}$ on $X_{1}\times X_{2}$ and $X_{2}\times X_{3}$, respectively. We denote by $\pi^{v}_{13}$ the marginal on $X_{1}\times X_{3}$. Similarly, the gluing lemma adapted to Radon measures and admissible couplings yields a measure $\pi^{e}$ on $\overline{Y_{1}}\times\overline{Y_{2}}\times\overline{Y_{3}}$ with marginals that agree with $\pi^{e}_{ij}$ when restricted to $\overline{Y_{i}}\times\overline{Y_{j}}\setminus\{(\partial_{Y_{i}},\partial_{Y_{j}})\}$ and that induces a zero cost on $(\partial_{Y_{i}},\partial_{Y_{j}})$. Then, we have
$d_{\mathrm{TpOT},p}(P_{1},P_{3})$
$\leq\left(\|d_{p}\circ(\iota_{1},\iota_{3})\|^{p}_{L^{p}(\pi^{e}_{13})}+\|k_{1}-k_{3}\|^{p}_{L^{p}(\pi^{v}_{13}\otimes\pi^{v}_{13})}+\|\omega_{1}-\omega_{3}\|^{p}_{L^{p}(\pi^{v}_{13}\otimes\pi^{e}_{13})}\right)^{1/p}$ (12)
@@
Here, we use $\|\cdot\|_{p}$ to indicate the standard $\ell^{p}$ norm on $\mathbb{R}^{3}$. The inequality (12) follows by sub-optimality, while (13) and (16) by marginal conditions. The triangle inequalities in $L^{p}$ and $d_{p}$ give (14), and the triangle inequality in $\ell^{p}$ gives (15). ∎
@@
Two measure topological networks $P=\left((X,k,\mu),(Y,\iota,\nu),\omega\right)$ and $P^{\prime}=\left((X^{\prime},k^{\prime},\mu^{\prime}),(Y^{\prime},\iota^{\prime},\nu^{\prime}),\omega^{\prime}\right)$ have distance $d_{\mathrm{TpOT},p}(P,P^{\prime})=0$ if and only if $P\sim_{w}P^{\prime}$.
@@
The “if” direction is clear, by the triangle inequality. Conversely, assume $d_{\mathrm{TpOT},p}(P,P^{\prime})=0$ and let $\pi^{v}$ and $\pi^{e}$ be couplings realising this distance (Proposition 1).
@@ 3.3 A topology-driven metric on point clouds
We now summarise the metric described above in the discrete setting of point clouds with finite number of points. Consider a point cloud $X=\{x_{1},\cdots,x_{N}\}\subseteq\mathbb{R}^{n}$. Any choice of filtration $\mathcal{K}$ over $X$ yields a corresponding persistence diagram $D=D(X,\mathcal{K})$. Choosing a representative cycle for each homology class in $D$ yields the set of generators $g$, and the corresponding PH-hypergraph $H=H(X,\mathcal{K},g)$. We can now endow $X$ with a measure topological hypergraph structure by considering
@@
Given two point clouds $X=\{x_{1},\cdots,x_{N}\}\subseteq\mathbb{R}^{n}$ and $X^{\prime}=\{x^{\prime}_{1},\cdots,x^{\prime}_{N^{\prime}}\}\subseteq\mathbb{R}^{n^{\prime}}$ endowed with filtrations, let $P_{X}$ and $P_{X^{\prime}}$ be their associated topological networks. In this concrete setting, we focus on the $p=2$ version of the metric, since it is the case where computation of terms quadratic in the couplings can be done efficiently [47, Proposition 1]. As was described in Remark 3, it is useful in applications to include tunable weights on each of the terms of TpOT. Using notation similar to that of [11], the distance $d_{\mathrm{TpOT},2}=d_{\mathrm{TpOT},\alpha,\beta}$ can be reformulated as
$\begin{split}d_{\mathrm{TpOT},\alpha,\beta}(P_{X},P_{X^{\prime}})=\min_{\pi^{v}\in\Pi(\mu,\mu^{\prime}),\>\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\quad\Biggl{(}&\alpha\left\langle L(C,C^{\prime}),\pi^{v}\otimes\pi^{v}\right\rangle\\ &+(1-\alpha)\left\langle\widetilde{C}(D,D^{\prime}),\pi^{e}\right\rangle\\ &+\beta\left\langle L(\omega_{H},\omega_{H^{\prime}}),\pi^{v}\otimes\pi^{e}\right\rangle\Biggr{)}^{1/2}.\end{split}$ (17)
@@
  • All inner products, denoted by angle brackets, in this expression are Frobenius inner products of matrices.
  • $C\in\mathbb{R}^{|X|\times|X|},C^{\prime}\in\mathbb{R}^{|X^{\prime}|\times|X^{\prime}|}$ are pairwise affinity matrices for points in $X$ and $X^{\prime}$ respectively, and $D,D^{\prime}$ are the persistence diagrams associated to the given filtrations $\mathcal{K}(X)$ and $\mathcal{K}(X^{\prime})$.
  • $\widetilde{C}(D,D^{\prime})\in\mathbb{R}^{(|D|+1)\times(|D^{\prime}|+1)}$ denotes the augmented squared Euclidean cost matrix for persistence diagrams, whose last column and row correspond to transport to and from $\partial_{\Lambda}$ respectively, following the diagonal projection $\pi_{\partial_{\Lambda}}$ (see [37, Equation 8]).
  • $L(C,C^{\prime})\in\mathbb{R}^{|X|\times|X^{\prime}|\times|X|\times|X^{\prime}|}$ corresponds to the coefficients of the Gromov-Wasserstein distortion functional, given by $L(C,C^{\prime})_{ijk\ell}=\frac{1}{2}|C_{ik}-C^{\prime}_{j\ell}|^{2}$.
  • $\omega_{H}\in\mathbb{R}^{|X|\times(|D|+1)}$ and $\omega_{H^{\prime}}\in\mathbb{R}^{|X^{\prime}|\times(|D^{\prime}|+1)}$ the hypernetwork functions of the PH-hypergraphs $H$ and $H^{\prime}$, respectively. In particular, we include an additional last column of zeros representing the diagonal $\partial_{\Lambda}$.
  • $L(\omega_{H},\omega_{H^{\prime}})$ corresponds to the coefficients of the co-optimal transport distortion functional, and is given by:
    $L(\omega,\omega^{\prime})_{ijkl}=\frac{1}{2}\begin{cases}|\omega_{ik}-\omega^{\prime}_{jl}|^{2},&\text{ for }1\leq k\leq|D|,1\leq l\leq|D^{\prime}|,\\ |\omega_{ik}|^{2},&\text{ for }1\leq k\leq|D|,l=|D^{\prime}|+1\\ |\omega^{\prime}_{jl}|^{2},&\text{ for }k=|D|+1,1\leq l\leq|D^{\prime}|\\ 0,&\text{ for }k=|D|+1,l=|D^{\prime}|+1.\end{cases}$ (18)
    In other words, the cost of matching any “real” (i.e., not the virtual one introduced at $|D|+1$) edge to the diagonal (representing $\partial_{\Lambda}$) is given by the squared $L^{2}$-norm of the edge, while matching the diagonal with the diagonal has zero cost.
@@
The solution of the TpOT problem in (17) involves determining an optimal pair of couplings $(\pi^{v},\pi^{e})$. The coupling $\pi^{v}$ induces a transport plan between points in the point clouds $X,X^{\prime}$. When $\beta>0$ and $10<\alpha<$ 1, this transport balances between preserving topological features and pairwise distances. The coupling $\pi^{e}$ induces a partial matching between homology classes in the persistence diagrams $D,D^{\prime}$. For $10<\alpha<$ 1, this matching is informed by the proximity of points, corresponding to representative cycles of classes in $D,D^{\prime}$ in the persistence diagram space $\Lambda$. Implementation details are described in Section 5, where we also provide numerical examples.
@@ 3.4 Numerical algorithms
We now aim at numerical algorithms for approximating the solution to the TpOT problem in practice. Starting from (17), we consider a entropically regularised variant of the TpOT problem by adding an entropic regularisation to the transport plans $(\pi^{e},\pi^{v})$. Writing
$\mathsf{L}(\pi^{v},\pi^{e}):=\alpha\langle L(C,C^{\prime}),\pi^{v}\otimes\pi^{v}\rangle+(1-\alpha)\langle\widetilde{C}(D,D^{\prime}),\pi^{e}\rangle+\beta\langle L(\omega_{H},\omega_{H^{\prime}}),\pi^{v}\otimes\pi^{e}\rangle,$ (19)
@@
$\min_{\pi^{v}\in\Pi(\mu,\mu^{\prime}),\>\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}$ $\mathsf{L}(\pi^{v},\pi^{e})+\varepsilon_{v}\operatorname{KL}(\pi^{v}|\mu\otimes\mu^{\prime})+\varepsilon_{e}\operatorname{KL}(\pi^{e}|\widetilde{\nu}\otimes\widetilde{\nu}^{\prime}),$ (20)
where $\varepsilon_{v},\varepsilon_{e}>0$ specify the strength of the entropic regularisation, and allow us to utilise fast, smooth optimisation techniques [49, 46, 47]. In particular, we can find a local minimum by projected gradient descent [47, Section 2.3], where both the gradient and the projection are calculated with respect to the KL-divergence. This leads to the following iterative scheme:
$\begin{split}\pi^{v}_{t+1}&\leftarrow\operatorname{Proj}_{\Pi(\mu,\mu^{\prime})}^{\operatorname{KL}}\left[e^{-\varepsilon_{v}^{-1}\nabla_{v}\mathsf{L}(\pi^{v}_{t},\pi^{e}_{t})}(\mu\otimes\mu^{\prime})\right],\\ \pi^{e}_{t+1}&\leftarrow\operatorname{Proj}_{\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}^{\operatorname{KL}}\left[e^{-\varepsilon_{e}^{-1}\nabla_{e}\mathsf{L}(\pi^{v}_{t},\pi^{e}_{t})}(\widetilde{\nu}\otimes\widetilde{\nu}^{\prime})\right].\end{split}$ (21)
Each of these projections can be calculated by matrix scaling using the Sinkhorn algorithm [46, Chapter 4], specifically:
$\operatorname{Proj}^{\operatorname{KL}}_{\Pi(\alpha,\alpha^{\prime})}\left(e^{\frac{-C}{\varepsilon}}(\alpha\otimes\alpha^{\prime})\right)=\min_{\pi\in\Pi(\alpha,\alpha^{\prime})}\langle C,\pi\rangle+\varepsilon\operatorname{KL}(\pi|\alpha\otimes\alpha^{\prime}).$ (22)
The gradients of $\mathsf{L}$ in $(\pi^{v},\pi^{e})$ have the following closed-form expressions:
$\nabla_{v}\mathsf{L}(\pi^{v},\pi^{e})$ $=2\alpha L(C,C^{\prime})\otimes\pi^{v}+\beta L(\omega_{H},\omega_{H}^{\prime})\otimes\pi^{e},$ (23)
$\nabla_{e}\mathsf{L}(\pi^{v},\pi^{e})$ $=(1-\alpha)\widetilde{C}[D,D^{\prime}]+\beta L(\omega_{H}^{\top},{\omega_{H}^{\prime}}^{\top})\otimes\pi^{v}.$ (24)
In the unregularised case (i.e., when $\varepsilon_{v}=\varepsilon_{e}=0$), an alternating minimisation in $(\pi^{v},\pi^{e})$ can be used to find a local minimum. Fixing $\pi^{e}$, the minimisation problem in $\pi^{v}\in\Pi(\mu,\mu^{\prime})$ is
@@
This falls within the fused Gromov-Wasserstein framework which was introduced and studied in detail by [59, 58]. A local minimum can be found using a conditional gradient method [59, Algorithm 1]. On the other hand, fixing $\pi^{v}$ and minimising in $\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})$, we have
$\min_{\pi^{e}\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\>\langle M,\pi^{e}\rangle,\quad\text{ where }M=\pi^{v}\otimes L(\omega_{H},\omega_{H^{\prime}})+(1-\alpha)\widetilde{C}(D,D^{\prime})$ (26)
This amounts to a standard optimal transport problem, and we have the identity (see [47, Proposition 1])
@@
For clarity, we provide pseudocode below for the entropy regularised approach (Algorithm 1) as well as the unregularised approach (Algorithm 2).
Algorithm 1 Entropy-regularised TpOT: Projected gradient descent
1:Input: Pairwise affinity matrices $C,C^{\prime}$, persistence diagrams $D,D^{\prime}$, incidence matrices $\omega,\omega^{\prime}$, probability distributions $\mu,\mu^{\prime}$, positive measures $\nu,\nu^{\prime}$
2:Parameters: Weights $\alpha\in[0,1],\beta\geq 0$, entropic regularisation coefficients $\varepsilon_{v},\varepsilon_{e}>0$.
3:Initialise couplings: $\pi^{v}_{1}\leftarrow\mu\otimes\mu^{\prime},\>\pi^{e}_{1}\leftarrow\widetilde{\nu}\otimes\widetilde{\nu}^{\prime}.$
[…]
5:    $M^{v}\leftarrow\nabla_{v}\mathsf{L}(\pi_{t}^{v},\pi_{t}^{e})$ using (23)
6:    $\pi^{v}_{t+1}\leftarrow\min_{\pi\in\Pi(\mu,\mu^{\prime})}\>\langle M^{v},\pi\rangle+\varepsilon_{v}\operatorname{KL}(\pi|\mu\otimes\mu^{\prime})$
7:    $M^{e}\leftarrow\nabla_{e}\mathsf{L}(\pi_{t}^{v},\pi_{t}^{e})$ using (24)
8:    $\pi^{e}_{t+1}\leftarrow\min_{\xi\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\>\langle M^{e},\xi\rangle+\varepsilon_{v}\operatorname{KL}(\xi|\widetilde{\nu}\otimes\widetilde{\nu}^{\prime})$
9:end for
10:Output: couplings $(\pi^{v},\pi^{e})$
Algorithm 2 Unregularised TpOT: Block-coordinate descent
1:Input: Pairwise affinity matrices $C,C^{\prime}$, persistence diagrams $D,D^{\prime}$, incidence matrices $\omega,\omega^{\prime}$, probability distributions $\mu,\mu^{\prime}$, positive measures $\nu,\nu^{\prime}$
2:Parameters: Weights $\alpha\in[0,1],\beta\geq 0$.
3:Initialise couplings: $\pi^{v}_{1}\leftarrow\mu\otimes\mu^{\prime},\>\pi^{e}_{1}\leftarrow\widetilde{\nu}\otimes\widetilde{\nu}^{\prime}.$
[…]
5:    $\pi^{v}_{t+1}\leftarrow\min_{\pi\in\Pi(\mu,\mu^{\prime})}\alpha\langle L(C,C^{\prime}),\pi\otimes\pi\rangle+\beta\langle L(\omega_{H},\omega_{H^{\prime}})\otimes\pi^{e}_{t},\pi\rangle$
6:    $M^{e}\leftarrow\pi^{v}_{t+1}\otimes L(\omega_{H},\omega_{H^{\prime}})+(1-\alpha)\widetilde{C}(D,D^{\prime})$
7:    $\pi^{e}_{t+1}\leftarrow\min_{\xi\in\Pi(\widetilde{\nu},\widetilde{\nu}^{\prime})}\>\langle M^{e},\xi\rangle$
8:end for
9:Output: couplings $(\pi^{v},\pi^{e})$
Note that the standard entropic regularisation term $\text{KL}(\pi_{e}|\widetilde{\nu}\otimes\widetilde{\nu}^{\prime})$ yields inhomogeneity (see for instance [36]), and thus, the entropic term may dominate when the cardinality of the persistent diagram increases. While we believe that the approach used in this preliminary work is sufficient for the current analysis, this effect is worth investigating further and could be an important focus for future research.